OpenStudy (anonymous):

Got confused by this limit at infinity problem I do know that the answer is -9/2 though I don't get much on how to get that answer I know how to do it with fractions though. so lim x -> +infinity (sqrt(x^4-9x^2)-x^2)

OpenStudy (anonymous):

\[\lim_{x \rightarrow +\infty} (\sqrt[4]{x^4-9x^2} - x^2)\] better visual lol

OpenStudy (mathmale):

I'd suggest that you multiply (and then divide) this expression by its conjugate. That way you'll end up with a more familiar form: a fraction. Try it, please.

OpenStudy (anonymous):

i got \[\frac{ -9x^2}{ \sqrt[4]{x^4-9x^2}+x^2 } \] i am going the right way? XD

OpenStudy (anonymous):

I probably need some sleep now if ever you come back could you please tell the next steps I'll get back to it in the morning lol.

OpenStudy (mathmale):

Think of what would happen to this expression if x were now to increase without bound. Note that x^4 is >>>x^2 in this scenario. does a limit seem any more tangible now?